OpenStudy (anonymous):

Solve for cos^2x=3/4

OpenStudy (anonymous):

but that's for cos^2x and it's equal to 3/4..how would you do that ?

OpenStudy (anonymous):

cos^2(x) is the same as saying (cosx)^2

OpenStudy (anonymous):

how do we account for the 3/4 part?

OpenStudy (anonymous):

(cosx)^2=3/4 cosx=sqrt(3/4) cosx=sqrt(3)/2 x=cos-1(0.866025) x=0.523599 Correct me if im wrong. Do you know the answer?

OpenStudy (anonymous):

no unfortunately I don't have the answer but thank you for helping and I'll try that.

OpenStudy (anonymous):

Trust me there is no other way.

OpenStudy (anonymous):

Thanks a lot for your help!

OpenStudy (anonymous):

i think 30 degree is the ans